what is the heat capacity of 100 kg of a coolant that has a specific heat capacity of 2420 J/KgK


1
Expert's answer
2022-05-23T10:17:44-0400

The heat capacity

"C=cm\\
=\rm 2420\: J/(K\cdot kg)*100\: kg=2.42*10^5\: J/K"